# Learning Object Placement Programs for Indoor Scene Synthesis with Iterative Self Training

### v1

@@ Abstract.
Data driven and autoregressive indoor scene synthesis systems generate indoor scenes automatically by suggesting and then placing objects one at a time. Empirical observations show that current systems tend to produce incomplete next object location distributions. We introduce a system which addresses this problem. We design a Domain Specific Language (DSL) that specifies functional constraints. Programs from our language take as input a partial scene and object to place. Upon execution they predict possible object placements. We design a generative model which writes these programs automatically. Available 3D scene datasets do not contain programs to train on, so we build upon previous work in unsupervised program induction to introduce a new program bootstrapping algorithm. In order to quantify our empirical observations we introduce a new evaluation procedure which captures how well a system models per-object location distributions. We ask human annotators to label all the possible places an object can go in a scene and show that our system produces per-object location distributions more consistent with human annotators. Our system also generates indoor scenes of comparable quality to previous systems and while previous systems degrade in performance when training data is sparse, our system does not degrade to the same degree.
Indoor Scene Synthesis, Program Induction, Domain Specific Language, Transformers,
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: TOGjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Scene understandingccs: Computing methodologies Unsupervised learningccs: Computing methodologies Neural networksccs: Computing methodologies Statistical relational learning
@@
Figure 1. Autoregressive indoor scene synthesis systems perform scene synthesis by iteratively placing objects. Given a partial scene and an object to add, they predict possible places that object can go. We visualize the placement options produced by our and previous systems for the inputs shown on the left. Note how previous systems predict incomplete distributions while our system produces a variety of placement options.
[…]
TODO: teaser description
1. Introduction
Humans spend a majority of their time indoors, and our media and technical applications reflect this lived experience. Augmented / virtual reality (AR/VR), movies, and video games all rely on virtual indoor spaces. Computer vision and robotics researchers working on tasks such as scene understanding and robotic navigation also rely on 3D scenes (Gadre et al., 2022; Puig et al., 2023). Manually authoring indoor scenes however is laborious, time consuming, and often requires expert knowledge — a process which does not scale.
Indoor scene synthesis systems which can automatically generate indoor scenes have emerged as a solution to these demands. While these methods take on a variety of formulations, we focus on data-driven and autoregressive systems (Paschalidou et al., 2021; Ritchie et al., 2018). Data-driven methods learn to generate scenes similar to those in an example dataset. This formulation is useful in instances where a user wants to generate scenes that mimic design patterns present in a set of exemplars. Autoregressive systems perform scene synthesis by either iteratively placing objects in an empty scene until it is complete, or scene completion from a given partial scene via the same methodology. For applications in interactive scene generation (5D, 2011; RoomSketcher, 2024), these systems can suggest both the next object to place and where in the scene it should be placed.
@@ 1. Introduction
Empirical observations of next object location distributions predicted by data-driven autoregressive systems reveal their tendency to become incomplete with respect to human-authored placement rules. For example, a distribution which only places a bed in the corner of an empty room is incomplete since one could place it along any of the walls. This flaw both limits the diversity of possible final scenes for scene synthesis and completion tasks, and limits their overall usefulness in an interactive application. Data-driven systems overfit to particular object placements seen during training. When 3D scene data is sparse, this effect is amplified as there are not enough examples to learn placement rules effectively.
@@
Standard solutions to reducing overfitting in scene synthesis systems are unreliable and come at a cost. Scaling the amount of scene data a system trains on can help an autoregressive system learn a more complete distribution, but this solution is expensive. Neural networks are also known to represent the most common inputs over the rare (Che et al., 2017) so these models might still miss placement modes. Model regularization and stopping training early is another option, but in practice it is hard to balance mode coverage and specificity. Better mode coverage often comes at the cost of decreased final scene quality.
@@
Humans craft their indoor environments to reflect their needs, desires, and tastes — a tv stand in front of a couch, a nightstand within arms reach of a bed, a wardrobe against a wall. Object-to-object and object-to-room relationships dictate the rules which guide object arrangements. A neural-network based system encodes these rules implicitly making them hard to control and align with our intuitions. Symbolic constructs such as programs can succinctly represent these rules. Their structured representation makes it easier to incorporate prior knowledge, edit the rules they represent, and ensure desirable characteristics such as continuity and completeness upon execution. We hypothesize that learning a symbolic representation, such as a Domain Specific Language (DSL) for specifying functional constraints can produce more complete per-object location distributions.
@@
We propose a new approach to indoor scene synthesis. Our system is autoregressive, placing objects one at a time. Rather than task a generative model with predicting possible object placements however, we instead ask it to predict a DSL program. We design this DSL with functional constraints that explicitly represent human activity and inter-object relationships. Programs from this DSL are relational layout programs which, given a partial scene and object to place, produce a mask representing all the possible object locations. We incorporate this language into a learning-based framework to learn how to automatically infer programs from partial scenes and query objects. Our system addresses the problem of incomplete next object location distributions that previous systems suffer from.
@@
We use a transformer-based generative model to generate programs automatically. Available 3D scene data contain no “ground truth” programs which can supervise said model so we introduce an iterative self-training scheme inspired by prior work in unsupervised visual program inference (Jones et al., 2022) that improves the performance of our system.
@@
To quantify the empirical observations mentioned previously we also introduce a new evaluation procedure which captures how well a system models per-object location distributions. We ask human annotators to label all the possible places an object can go in a scene and compare this rule against one produced by the system in question. We show how our system produces per-object location distributions that are more consistent with human annotators. While other systems show consistent degradation in per-object location modeling with less scene data, our system does not degrade in performance to the same degree.
In summary our contributions are:
@@
  • A new approach to indoor scene synthesis where we predict a relational layout program from a given partial scene and object to place, and execute that program to predict possible object placement locations
  • A bootstrapped self-training algorithm that improves our system’s performance
@@
2. Related Work
Indoor Scene Synthesis
@@ 2. Related Work
Before the existence of large indoor scene datasets (Fu et al., 2021) and 3D deep learning algorithms, researchers approached indoor scene synthesis with explicit rules such as statistical relationships between objects (Yu et al., 2011), programmatically-defined constraints  (Yeh et al., 2012), design principles (Merrell et al., 2011), or heuristics for human activity (Fu et al., 2017; Fisher et al., 2015). Similarly, our DSL also encodes object relationships and human activity explicitly.
@@
Deep learning methods enabled end-to-end and differentiable systems to learn scene piors automatically from large scene datasets. A scene graph is a popular representation with works using a graph neural network (Wang et al., 2019; Gao et al., 2023), recursive neural network (Li et al., 2018) or diffusion model (Tang et al., 2024) to learn these priors. Another line of work takes an image based approach and operates over the top down view of the scene with CNNs (Ritchie et al., 2018; Wang et al., 2018). Recent transformer based approaches (Paschalidou et al., 2021; Para et al., 2020; Wang et al., 2020) have also found success working directly with the 3D bounding box information of objects in the scene. Our system leverages recent advances in 3D deep learning, using the transformer architecture as the backbone for scene generation. Instead of having the network directly predict object placements however, our generative model writes programs defined by our DSL.
Another line of work is zero-shot indoor scene synthesis. These methods do not require training on 3D scene datasets and instead rely on latent scene knowledge embedded within Large Language Models (LLMs) to generate scenes. Holodeck (Yang et al., 2023) generates 3D environments for embodied AI training by sampling spatial relational constraints from an LLM and then optimizing the layout with a physical solver. Another work designs a LLM-based program synthesizer whose programs are specified by a DSL (Aguina-Kang et al., 2024). This DSL is a scene description program that specifies the objects and their spatial layout. We also generate DSL programs whose primitives are relational constraints, but our programs specify possible object placements rather than an entire scene. Our work also requires 3D scene data to train on and aims to match the distribution of input scenes rather than generate any indoor scene zero-shot.
Visual Program Inference
@@
Visual Program Inference (VPI) aims to automatically infer programs that explain visual data (Ritchie et al., 2023). If the visual data of interest comes with ground truth programs, supervised learning is an obvious option (Wu et al., 2021; Willis et al., 2021; Xu et al., 2022). In most domains however, programs for visual data are not readily accessible. Unsupervised learning, and in particular bootstrapping (Ellis et al., 2021; Liang et al., 2016), is one option for extracting and improving programs from visual data.
@@
Bootstrapping methods search for “good” programs, retrain on these programs, and then repeat. PLAD (Jones et al., 2022) groups different bootstrapping methods under a single conceptual framework and applies it to VPI on 3D and 2D shapes. The method sources programs from both a recognition model that matches input shapes with programs as “pseudo-labels” and pairs of synthetic shapes and their programs for an approximation of the target distribution. Later works works which built on the PLAD framework (Ganeshan et al., 2023; Jones et al., 2024) searched for new programs by editing existing ones. SIRI (Ganeshan et al., 2023) used domain specific operations to edit a subset of programs at a time for distributional stability. Our algorithm is an instance of PLAD and like SIRI we use domain-specific editing operations.
Figure 2. Inference Pipeline: Our system is autoregressive, placing objects one at a time. Given a (1) partial scene and query object we (2) sample a DSL program from a generative model. (3) Upon execution, this produces a binary mask of possible centroid locations for the query object in the scene. (4) Sampling this mask produces valid object locations of the query object.
[…]
TODO
@@
Figure 3. We develop a program bootstrapping algorithm which discovers programs automatically from scene data. (1) We start by extracting programs with geometric heuristics and then training our (2) model on these initial programs. (3) We propose new programs by deleting constraints from both the inferred and original programs. (4) These candidate programs are then filtered with a scene real fake classifier to remove ”bad” programs. (5) Domain specific operations combine ”good” candidate programs together, and insert them back into the training set.
[…]
TODO
3. Method Overview
@@
The following steps constitute our full system. Figure 2 shows our inference pipeline and Figure 3 shows how we train our system.
Defining a DSL to describe object placement distributions
@@
To represent object placement distributions in semantically meaningful terms and align placement rules closer to ones a human might write, we introduce a domain specific language (DSL). Section 4 describes this language and its motivations in more depth.
Automatic scene synthesis with programs
@@
To perform scene completion and scene synthesis we design a generative model which generates programs automatically. Section 5 describes this model.
Program bootstrapping
@@
Current indoor scene datasets do not contain programs to train on, so we introduce a program bootstrapping algorithm to discover these programs and boost system performance. Section 6 describes this algorithm.
4. Language Design
@@
Figure 4. Program Example: Given a partial scene and object to add, our DSL program outputs a binary mask representing possible placements of that object. Programs take on the structure of Constructive Solid Geometry (CSG) trees where each leaf node is a functional constraint that describes object function. Upon execution, these constraints produce binary masks which are combined according to the structure of the tree.
[…]
TODO
@@
Our DSL programs take as input a partial scene and the next object to place. They then output a binary mask representing possible centroid locations of the query object. These programs take on the same structure of a Constructive Solid Geometry (CSG) tree, but instead of operating over a continuous 3D space, the programs operate over 2D masks. Leaf nodes of this tree are functional constraints that explicitly represent human activity and inter-object relationships. When executed these constraints produce binary masks.
@@
Binary masks are discretized along 3 dimensions. The three dimensions represent the width of the room, the height of the room, and the possible orientations of the object to place respectively. This third dimension of the mask is necessary because the validity of an object’s centroid position is dependent on its orientation. The orientation of an object dictates the centroid locations the program executor predicts. We represent the orientation of an object as its rotation about the up axis of the room and snap it to one of the cardinal directions (N, E, S, W). Most objects in 3DFRONT (Fu et al., 2021) are axis aligned, so our language can model most object placements. Incorporating additional degrees of freedom for more complex scenarios is an avenue for future work.
Figure 4 shows an example program, its inputs, and its outputs.
4.1. Design Goals
We hypothesize that forcing our system to explain object placements with logical rules and semantically meaningful terms will produce placement rules more consistent with human intuition. Humans would likely describe object placement rules in terms of higher order concepts such as alignment, attachment, or reachability. They might also then use boolean statements to specify how those concepts should interact with each other. For example, one might describe a nightstand as both being attached to the side of a bed and aligned in its forward facing direction. A wardrobe could be attached to the left wall or the right wall. Programs excel at representing higher-order concepts and CSG is a straightforward choice for translating boolean statements to the visual domain. We are also inspired by previous work (Sharma et al., 2018; Ellis et al., 2019) in inferring and editing CSG programs from input geometry as we want a representation conducive to visual program induction.
4.2. Constraint Specification
@@
Functional constraints fall under two categories. Location constraints (attach and reachable_by_arm) assume nothing about an object’s orientation and predict an object’s possible centroid locations. Orientation constraints (align and face) constrain the possible orientations of the object within a scene in addition to the location.
@@
5 total directions are specified in the language (Up, Down, Left, Right, Null) and all directions are specified within the local coordinate frame of the reference object. The Null direction exists because orientation constraints do not need a direction specification.
Visualizations of each constraint alongside further implementation details are shown in the supplemental materials.
@@
5. Generating programs
Figure 5. Our generative model takes as input a partial scene and object and predicts a program written in our DSL. We formulate this modeling task as a seq-2-seq problem. (1) We first vectorize and then embed both the input objects and program. The structure of the program tree and the constraint attributes are embedded as seperate sequences. (2) Our first transformer encoder decoder pair predicts the structure of the program from the input objects. (3) Our second transformer encoder decoder pair predicts the constraint attributes from the object and structure embeddings.
[…]
TODO
The design of our generative model takes inspiration from previous work which use partial program intermediates to aid in program synthesis (Solar-Lezama, 2008; Ellis et al., 2021). These works use a partial programs to express the high-level structure, but leaves holes for low level implementation details. In similar fashion, we predict the high level topological structure of our program first with one network, and then ask a second network to fill in the low level instantiation details. This approach helps reduce overfitting to particular structure and parameter combinations.
5.1. Overview
We treat program synthesis as a sequence to sequence (seq-2-seq) translation task. The input sequence are the objects in the room and an object to place. The output or target sequence is the program. We train two transformer (Vaswani et al., 2017) encoder-decoder pairs in a two-pass approach. The first encoder-decoder pair takes in object encodings and outputs the structure or topology of the target program. The second encoder-decoder pair takes as input the source and target sequences of the first pass concatenated together and outputs the attributes of all the program’s constraints as a flattened list. Figure 5 shows the network architecture.
5.2. Object Encoding

### v2

# Learning to Place Objects with Programs and Iterative Self Training

@@ Abstract
In this work we study indoor scene object placement. Given a 3D indoor scene and an object, the task is to predict placement locations within the scene. Empirical observations of data-driven approaches to the problem show their tendency to miss placement modes. We introduce a system which helps to address this flaw. We design a Domain Specific Language (DSL) that specifies object relational constraints. Upon execution, programs from our language predict possible placements from a partial scene and object. We design a generative model which writes these programs automatically. Available 3D scene datasets do not contain programs to train on, and naively extracted programs only predict the original placement location of scene objects. Training on these programs results in subpar performance so we introduce a new program bootstrapping algorithm that improves our system’s performance compared to the naive approach. To quantify our qualitative observations, we introduce a new evaluation procedure which captures how well a system models per-object location distributions. We ask human annotators to label all the possible places an object can go in a scene and compare this set against locations produced by the system in question. Our system produces per-object location distributions more consistent with human annotators than those produced by existing data-driven approaches and a zero-shot approach using an LLM. While other systems degrade in performance when training data is sparse, our system does not degrade to the same degree.
@@
Figure 1: Indoor scene object placement is the task of predicting possible placement locations from a given 3D scene and object. We visualize the placement options produced by our and previous methods for the inputs shown on the left. Note how previous methods predict incomplete distributions while ours produces a variety of placement options.
1 Introduction
Placing objects in an indoor environment is a natural consequence of spending time in one. A variety of applications rely on predicting possible object locations from a given 3D scene and object. Computer vision and robotics researchers use this information in settings such as scene understanding and robot interaction  [10, 27, 28]. Users of augmented / virtual reality (AR/VR) and interactive scene generation applications  [1, 31] might want to automatically visualize placement options of an object in their indoor space. Predicting possible object locations is also a fundamental subroutine of several scene synthesis systems  [26, 30].
@@ 1 Introduction
Data-driven approaches to this problem  [26, 30] train on indoor scene datasets to learn placement distributions. Empirical observations of object location distributions predicted by these systems reveal their tendency to be incomplete, i.e they omit many plausible locations. For example, a distribution which only places a bed in the corner of an empty room is incomplete, since one could place it along any of the walls. These prior systems overfit to particular object placements seen during training, which limits their overall usefulness for the above applications.
@@
Standard solutions to reducing overfitting in these systems are unreliable and come at a cost. Scaling the amount of scene data a system trains on can help it learn a more complete distribution, but this solution is expensive. Neural networks are also known to represent the most common inputs over the rare [3], so these models might still miss placement modes. Model regularization and stopping training early is another option, but in practice it is hard to balance mode coverage with specificity.
@@
People use object-to-object and object-to-room relationships to guide object arrangement. Existing neural-network-based systems attempt to encode these rules implicitly, making them hard to control and align with our intuitions. On the other hand, symbolic representations can succinctly represent these rules. Their structured representation makes it easier to incorporate prior knowledge, edit the rules they represent, and ensure desirable characteristics such as completeness upon execution. We hypothesize that learning to produce a symbolic representation, such as programs in a Domain Specific Language (DSL), will result in more complete per-object location distributions.
@@
We propose a new approach to indoor scene object placement. Rather than tasking a generative model with predicting possible object placements, we instead ask it to predict a program in a DSL. Programs in this DSL are relational layout programs which explicitly represent human activity and inter-object relationships. Given a partial scene and object to place, they produce a binary mask representing all possible object locations. We incorporate this language into a learning-based framework to learn how to automatically produce programs from partial scenes and query objects. Our system helps to address the problem of incomplete next object location distributions that plague previous systems.
@@
We use a transformer-based generative model to produce programs. Available 3D scene dataset contain no “ground truth” location programs which can supervise the model. Programs extracted naively with geometric heuristics only predict the original placement location of scene objects. Training on these programs results in subpar performance, so we introduce an iterative self-training scheme which applies the PLAD framework  [18] — prior work in unsupervised visual program inference — to this new domain of object location programs. Our self-training algorithm improves our system’s performance compared to the naive extraction approach.
The original PLAD method assumes access to ground-truth shapes to learn to predict shape programs. Our setting is much harder because the ground truth location distributions which our programs are trying to match are not available. Our method only has access to single location samples during training. Despite this challenge, our method reproduces a good portion of these placement distributions. For each existing program, our approach proposes new programs, filters out noisy suggestions, and then combines “good” candidate programs together. Iteratively repeating this process results in programs that predict a variety of placement locations.
@@
To quantify the performance of object location prediction methods, we also introduce a new evaluation procedure which captures how well a system models per-object location distributions. We ask human annotators to label all the possible places an object can go in a scene and compare this set against locations produced by the system in question. Our system produces per-object location distributions more consistent with human annotators than those produced by existing data-driven approaches and a zero-shot approach using an LLM. While other systems show consistent degradation in per-object location modeling with less scene data, our system does not degrade in performance to the same degree.
In summary, our contributions are:
@@
  • A new approach for placing objects in indoor scenes where we predict a relational layout program from a given partial scene and object to place, and execute that program to predict possible object placement locations
  • A new bootstrapped self-training algorithm that adds placement modes to naive single location programs by iteratively proposing, filtering, and then combining programs together
@@
2 Related Work
@@
Indoor scene synthesis and object placement. Before the existence of large indoor scene datasets [8] and 3D deep learning algorithms, researchers positioned objects in scenes with explicit rules such as statistical relationships between objects [49], programmatically-defined constraints  [48], design principles [23], or heuristics for human activity [9, 7]. Similarly, our DSL also encodes object relationships and human activity explicitly.
@@
Deep learning enabled a variety of approaches for learning scene priors from large datasets. A scene graph is a popular representation with works using a graph neural network [41, 12], recursive neural network [21] or diffusion model [37, 35, 16] to learn these priors. Image based approaches operate over the top down view of the scene with CNNs [30, 42]. Transformer based approaches [26, 25, 43] found success working directly with the 3D bounding box information of objects in the scene. Our system leverages recent advances in 3D deep learning, using the transformer architecture as the backbone for scene generation. Instead of having the network directly predict object placements however, our generative model writes programs defined by our DSL.
Other works source scene priors beyond information distilled from 3D scene datasets. Zero-shot approaches leverage the latent scene knowledge embedded within Large Language Models (LLMs) to position objects and generate scenes. LayoutGPT  [6] directly predicts the positions and orientations of objects. Predicting constraints and then solving them for possible placement locations is another popular approach  [47, 2, 34, 17]. We also generate and then solve constraint programs. Our method however relies solely on information within 3D scene datasets.
@@
Visual Program Inference. Visual Program Inference (VPI) aims to automatically infer programs that explain visual data [29]. If the visual data of interest comes with ground truth programs, supervised learning is an obvious option [45, 44, 46]. In most domains however, programs for visual data are not readily accessible. Unsupervised learning, and in particular bootstrapping [5, 22], is one option for extracting and improving programs from visual data.
@@
Bootstrapping methods search for “good” programs, retrain on these programs, and then repeat. PLAD [18] groups different bootstrapping methods under a single conceptual framework and applies it to VPI on 3D and 2D shapes. Later works which built on the PLAD framework [11, 19] searched for new programs by editing existing ones. SIRI [11] used domain specific operations to edit a subset of programs at a time for distributional stability. Our algorithm is an instance of PLAD, and like SIRI, we use domain-specific editing operations.
@@
Figure 2: Our self-training algorithm discovers programs automatically from scene data. (1) Naively extracted programs serve as initial training data for our (2) model. (3) Deleting constraints from the inferred and original programs produce candidate programs which are then filtered by a (4) classifier. (5) “Good” programs are combined with domain specific operations, and then inserted back into the training set.
3 Method Overview
@@
Figure 1 shows our inference pipeline and Figure 2 shows how we train our system. Our full system consists of the following steps:
@@
Defining a DSL to describe object placement distributions. To represent object placement distributions in semantically meaningful terms and align placement rules closer to ones a human might write, we introduce a domain specific language (DSL). Section 4 describes this language and its motivations in more depth.
@@
Learning to generate programs. We design a generative model which writes programs automatically from a partial scene and object to add. Section 5 describes this model.
@@
Program bootstrapping. Current indoor scene datasets do not contain programs to train on, so we introduce a program bootstrapping algorithm to discover these programs and boost system performance. Section 6 describes this algorithm.
4 Language Design
@@
Figure 3: Example Program: Given a partial scene and object to add, our DSL program outputs a binary mask representing possible placements of that object. Programs take on the structure of Constructive Solid Geometry (CSG) trees where each leaf node is a constraint that describes object function. Upon execution, these constraints produce binary masks which are combined according to the structure of the tree.
@@
Our DSL programs take as input a partial scene and the next object to place. They then output a binary mask representing possible centroid locations of the query object. These programs take on the same structure of a Constructive Solid Geometry (CSG) tree, but instead of operating over a continuous 3D space, the programs operate over 2D masks. Leaf nodes of this tree are functional constraints that explicitly represent human activity and inter-object relationships. When executed, these constraints produce binary masks.
@@
Our binary masks are discretized along 3 dimensions: the width of the room, the height of the room, and the possible orientations of the object to place, respectively. This third dimension of the mask is necessary because the validity of an object’s centroid position is dependent on its orientation. We represent the orientation of an object as its rotation about the up axis of the room and snap it to one of the cardinal directions (N, E, S, W). Only 6% of objects in 3D-FRONT  [8] deviate more than 10 degrees from one of the cardinal directions so our language can model most object placements. Figure 3 shows an example program, its inputs, and its outputs.
We hypothesize that forcing our system to explain object placements with logical rules and semantically meaningful terms will produce placement rules more consistent with human intuition. CSG is a straightforward choice for translating logical statements into the visual domain and previous work [32, 4] has shown how CSG programs are conducive to to visual program induction.
4.1 Constraint Specification
@@
Constraints in our DSL fall under two categories. Location constraints (attach and reachable_by_arm) predict an object’s possible centroid locations for every orientation. Orientation constraints (align and face) constrain both location and orientation.
@@
Five total directions are specified in the language (Up, Down, Left, Right, Null), and all directions are specified within the local coordinate frame of the reference object. To standardize how constraints in the language are represented for neural net processing, all constraints take the same number of arguments. Since orientation constraints do not require a directional argument, we use a special Null direction for their direction argument value.
@@
5 Generating programs
Figure 4: Our model predicts DSL programs from a partial scene and object. We formulate this task as a seq-2-seq problem. (1) We vectorize and then embed both the input objects and program. The structure of the program tree and the constraint attributes are embedded as separate sequences. (2) Our first transformer encoder-decoder pair predicts the structure of the program from the input objects. (3) Our second transformer encoder decoder pair predicts the constraint attributes from the object and structure embeddings.
Previous approaches to program synthesis use partial programs to express the high-level structure but leave holes for low level implementation details [33, 5]. In similar fashion, our model uses one network to predict the topological structure of our program, and then another to fill in the details.
5.1 Overview
We treat program synthesis as a sequence to sequence (seq-2-seq) translation task. The input sequence are the objects in the room and an object to place. The output sequence is the program. We train two transformer [39] encoder-decoder pairs. The first pair takes in object encodings and outputs the “structure” of the program. The second pair takes in both the object encodings and program structure and outputs the constraint attributes. Figure 4 shows the network architecture.
5.2 Object Encoding